### v1

# Low-Redundant Optimization for Large Language Model Alignment

Abstract
Large language models (LLMs) are still struggling in aligning with human preference in complex tasks and scenarios. They are prone to overfit into the unexpected patterns or superficial styles in the training data. We conduct an empirical study that only selects the top-10% most updated parameters in LLMs for alignment training, and see improvements in the convergence process and final performance. It indicates the existence of redundant neurons in LLMs for alignment training. To reduce its influence, we propose a low-redundant alignment method named ALLO, focusing on optimizing the most related neurons with the most useful supervised signals. Concretely, we first identify the neurons that are related to the human preference data by a gradient-based strategy, then identify the alignment-related key tokens by reward models for computing loss. Besides, we also decompose the alignment process into the forgetting and learning stages, where we first forget the tokens with unaligned knowledge and then learn aligned knowledge, by updating different ratios of neurons, respectively. Experimental results on 10 datasets have shown the effectiveness of ALLO. Our code and data are available at https://github.com/RUCAIBox/ALLO.
1 Introduction
Alignment with human preferences has become a desired property of LLMs Askell et al. (2021); Ouyang et al. (2022); Zhao et al. (2023), e.g., helpfulness, honesty, and harmlessness, and reinforcement learning from human feedback (RLHF) Christiano et al. (2017); Zheng et al. (2023) is a widely utilized technique for achieving it. Typically, RLHF aims to fine-tune LLMs on human preference data, to maximize and minimize the likelihood of generating the positive and negative responses, respectively. After RLHF training on corresponding datasets, LLMs can better follow user instructions Ouyang et al. (2022), solve complex problems Wang et al. (2023), and generate unbiased responses Bai et al. (2022a).
Figure 1: Training loss curve and benchmark performance of QA tasks using different trainable neurons in LLM. We perform alignment training using DPO Rafailov et al. (2023) on ECQA and QASC. The top/last-10% related neurons are selected based on the accumulated gradients during DPO training.
However, it is hard to train a well-aligned LLM for complex tasks and scenarios Feng et al. (2024); Gekhman et al. (2024). The key issue is that LLMs might overfit into the unexpected patterns or superficial styles in the human preference data (i.e., pairs of positive-negative response) Du et al. (2024). It is the side effect of their powerful learning capability derived from the large-scale trainable parameters Song et al. (2024); Meng et al. (2024a). Recently, a surge of work Frankle and Carbin (2019); Wang et al. (2024b) has found that each neuron (i.e., one of the trainable values of the parameter matrixes in LLMs) is relevant with special knowledge, and the neurons in LLMs are generally sparsely activated. Inspired by it, we consider whether the irrelevant neurons exist in LLMs during the alignment process and the full-parameter trained LLMs might lead to redundant updates on alignment-irrelevant neurons. Thus, we conduct the empirical experiment using DPO algorithm, where we only update the top/last-10% neurons according to their accumulated gradient values, which indicate the relevance between the current neuron and the downstream scenario Pruthi et al. (2020). As shown in Figure 1, with the top-10% trainable neurons, LLMs can converge faster and achieve better performance than optimizing all the neurons. It indicates the existence of alignment-irrelevant neurons and redundant updates in DPO training, affecting the convergence and final performance. Besides the redundant neurons, previous work Lin et al. (2024) has shown the existence of redundant tokens in training data.
To reduce the influence of redundant updates, we aim to prune the alignment training, to focus on optimizing the most related neurons with the most useful supervised signals. Concretely, we first identify the neurons that are related to the human preference data, based on the accumulation values of gradients. Second, we identify the key tokens about human preference, and only compute loss on them for optimizing the alignment-related neurons. In this way, we perform a low-redundant optimization for aligning LLMs with humans, which reduces the redundancy of learning irrelevant tokens and training irrelevant neurons, avoiding their effect on alignment performance.
However, the involved tokens and neurons are not always consistent in the alignment objectives, since the alignment training focuses on both removing the unaligned knowledge and learning the aligned one. Therefore, we decompose the alignment process into the forgetting and learning stages, and adapt the low-redundant optimization strategy on them. For the forgetting stage, relatively fewer neurons are trained by unlearning algorithm Zhang et al. (2024) to forget the unaligned knowledge, and we leverage a token-level reward model Chen et al. (2024b) to identify the unaligned tokens required to be focused. For the learning stage, we train more neurons using the DPO algorithm, and also utilize its reward score to select the key tokens.
In this work, we proposed an ALignment method with Low-Redundant Optimization (ALLO) to fine-tune LLMs. In ALLO, we first identify the most important neurons by training a reference model. Then, for the forgetting and learning stages, we utilize NPO and DPO algorithms with unlearning and learning losses on corresponding key tokens respectively, to optimize the identified important neurons. To comprehensively assess the effectiveness of ALLO, we conduct extensive experiments on three downstream scenarios, i.e., question answering, mathematical reasoning, and instruction following, totally 10 datasets. Experiment results show that ALLO mostly outperforms competitive human alignment methods (e.g., SFT Ouyang et al. (2022), DPO Rafailov et al. (2023), PPO Schulman et al. (2017)), achieving a 9.7% maximum relative improvement over vanilla DPO.
2 Related Work
We introduce the related work from the perspectives of large language models, LLMs alignment, and unlearning of LLMs.
Large Language Models. LLMs have shown remarkable performance on various tasks qwe (2024); Meta (2024); Javaheripi et al. (2023). Generally, the training process of LLMs includes three stages, i.e., pre-training, supervised fine-tuning (SFT), and alignment Ouyang et al. (2022); Touvron et al. (2023). In the training process, previous work has selected valuable data to train the LLMs via leveraging gradient Xia et al. (2024) or perplexity Lin et al. (2024); Xie et al. (2023), Besides, synthetic training data from powerful LLMs (e.g., GPT-4, Claude 3) has been widely utilized for improving the weak LLMs Xu et al. (2023); Ben Allal et al. (2024); Liu et al. (2024), especially for specific scenarios (e.g., mathematical tasks or code synthesis tasks) Yue et al. (2023); Zhou et al. (2024). However, given the large expenses of the LLM training, existing work Hu et al. (2022); Li and Liang (2021); Dettmers et al. (2023) has revealed that training only a small number of the parameters can achieve comparable performance with whole-parameters training. In this work, we focus on the alignment stage and leverage the low-redundant optimization to improve the existing LLMs.
LLMs Alignment. RLHF is a critical algorithm of LLM alignment Christiano et al. (2017), usually leveraged to reduce hallucination Chaudhari et al. (2024) or further enhance the capacities of LLMs Chen et al. (2024b); Wang et al. (2023); Luo et al. (2023). Typically, a reward model will be trained on the preference data and leveraged to guide the reinforcement learning (RL) procedure Ouyang et al. (2022); Touvron et al. (2023); Zheng et al. (2023). Proximal policy optimization (PPO) has been widely adopted in RLHF Mnih et al. (2016); Zheng et al. (2023). Given the efficiency and expenses of the annotating process by human labeler, previous work has utilized the feedback from LLMs to instruct the RL process, named RLAIF Bai et al. (2022b); Yuan et al. (2024). Furthermore, to prevent the instability of RL, a series of work Park et al. (2024); Hong et al. (2024); Meng et al. (2024b) utilized a similar objective function with SFT to model human preference. Direct preference optimization (DPO) Rafailov et al. (2023) is representative work of non-RL alignment. In this work, we consider about how to unleash the potential of the non-RL method.
Unlearning of LLMs. Machine unlearning Cao and Yang (2015); Bourtoule et al. (2019); Wang et al. (2024a); Chen et al. (2024a) is an important technique for artificial intelligence systems to remove the knowledge about the restricted data (e.g., unauthorized books), while keeping other knowledge and abilities of the systems. To perform unlearning of LLMs, research has proposed several methods (e.g., Gradient Ascent Yao et al. (2023); Maini et al. (2024) and NPO Zhang et al. (2024)), directly training LLMs on the invalid dataset to make LLMs forget relative knowledge. Following the unlearning mechanism, in this work, we utilize an unlearning algorithm to correct the unaligned knowledge stored in the neurons of LLMs.
3 Preliminary
LLMs alignment refers to aligning the behaviors of LLMs to human preference, e.g., helpfulness, honesty, and harmlessness Askell et al. (2021). Existing work typically utilizes RLHF methods Christiano et al. (2017) to fine-tune LLMs using human preference data, for improving alignment. Formally, the human preference data is composed by input prompts, positive responses, and negative responses, denoted as $\mathcal{D}=\{\langle x_{i},y_{i}^{+},y_{i}^{-}\rangle\}_{i=1}^{n}$. The input prompt or response consists of a series of natural language tokens $\{t_{1},t_{2},\dots,t_{l}\}$. Given the input prompt $x$, we aim to train LLMs that tend to generate the well-aligned positive response $y^{+}$, while avoiding generating the unaligned negative one $y^{-}$. In this work, we focus on devising an effective training algorithm to improve the alignment of LLMs, which can be utilized to satisfy the diverse requirements in real world (e.g., instruction following and question answering).
According to our empirical study in Figure 1, updating only top-10% trainable neurons would achieve better performance than full-parameter tuning for alignment training. It indicates that there are redundant updates in the training process of LLMs, which may affect the alignment performance. To address it, in this work, we aim to perform parameter-efficient fine-tuning for reducing the redundant updates on unrelated neurons, to improve the alignment of LLMs. Given the training data, we first identify the highly-relevant neurons $\mathcal{N}=\{\theta_{i_{1}},\dots,\theta_{i_{k}}\}$ in the parameter matrices of LLMs, and perform low-redundant optimization on the LLM as:
$\theta_{i}^{t+1}=\begin{cases}\text{Optimizer}(\theta_{j}^{t},\nabla\theta_{j}^{t}),&\theta_{j}\in\mathcal{N}\\ \theta_{j}^{t},&\theta_{j}\notin\mathcal{N}\end{cases},$ (1)
where $\theta_{j}^{t}$ means the value of $j$-th neuron at the $t$-th step of training process, $\nabla\theta_{j}$ is the calculated gradient of $j$-th neuron for update.
Figure 2: The framework of our proposed alignment method ALLO. We first locate the key neurons in LLMs by computing the weight changes of the reference model.Then, based on the selected key neurons, we perform a fine-grained unlearning using NPO to help LLMs forget unaligned knowledge, and fine-grained learning using DPO to further align LLMs to human preference.
4 Approach
In this section, we introduce our proposed method ALLO, a low-redundant alignment method for fine-tuning LLMs. In ALLO, we compute loss on selected key tokens, and only optimize the selected important neurons. Concretely, we first train a reference model to locate the important neurons through gradient. Then, we identify the key tokens related to unaligned knowledge, and utilize the unlearning algorithm to update few neurons for forgetting them. Next, we leverage DPO algorithm to improve the alignment of the LLM, where the DPO reward is used for selecting the key tokens. The framework of ALLO is presented in Figure 2.
4.1 Locating Key Neurons
We compute the importance of all the neurons for the human preference data to locate the related key neurons. We first train a reference model on the given data using DPO algorithm, and then design an efficient approximate estimation of the neuron importance based on its updated weights.
Training Reference Model. We train the reference model on the human preference data, to obtain the updated values of all neurons for importance estimation. Thus, we select the same LLM as the backbone, and perform full-parameter fine-tuning using DPO algorithm on the entire dataset for one epoch. The training objective is:
$\mathcal{L}(d_{i})=-\log\sigma\left(\beta\log\frac{P(y_{i}^{+}|x_{i})}{P_{\text{ref}}(y_{i}^{+}|x_{i})}-\beta\log\frac{P(y_{i}^{-}|x_{i})}{P_{\text{ref}}(y_{i}^{-}|x_{i})}\right),$ (2)
where $\beta$ is a hyper-parameter, and $d_{i}=\langle x_{i},y_{i}^{+},y_{i}^{-}\rangle$ is a training instance. For the scenarios that only one human feedback is provided, we regard it as the positive one, and leverage the response generated from LLM as the negative one.
Neurons Importance Estimation. We aim to estimate the importance of each neuron for the given human preference dataset $\mathcal{D}$. As LLMs are generally trained by gradient descent algorithm, the gradient value of a training instance $d_{j}$ on the neuron $\theta_{i}$ can reflect its influence on the neuron Pruthi et al. (2020); Xia et al. (2024), denoted as:
$\text{Influence}(d_{j},\theta_{i})\propto\nabla_{\theta_{i}}\mathcal{L}(d_{j})$ (3)
For human alignment, we use the DPO training loss in Eq. 2 for influence estimation. In this way, we can accumulate the gradients for all the instances from the human preference dataset, to estimate the influence of the dataset on the neuron. Actually, the influence value also reflects the importance of the neuron for learning the dataset, as a large accumulated gradient value can denote more focus on training the neuron Pruthi et al. (2020). As we adopt the gradient descent algorithm, the gradients for all the instances have been computed and subtracted in the one-epoch training process. Thus, the difference between the neuron in the reference model $\theta^{\prime}_{i}$ and original model $\theta_{i}$ can be regarded as the approximate value of the estimated importance score:
$\mathcal{I}(\mathcal{D},\theta_{i})=\sum_{j=1}^{|\mathcal{D}|}\nabla_{\theta_{i}}\mathcal{L}(d_{j})\approx\frac{\theta^{\prime}_{i}-\theta_{i}}{\alpha},$ (4)
where $\alpha$ is the learning rate during DPO training. Based on the estimated importance score, we can rank all the neurons and select the most important ones for training.
4.2 Unaligned Knowledge Forgetting
For the forgetting stage, we utilize a token-level reward model that guides LLMs to focus on the tokens related to unaligned knowledge, and adopt a machine unlearning algorithm, i.e., NPO Zhang et al. (2024) that learns to forget them.
Unalignment-Related Tokens Identification. We train a token-level reward model to score tokens in the negative responses, according to their effect on unalignment. Following existing work Chen et al. (2024b), we distill the capability of a strong LLM (i.e., GPT-4 OpenAI (2023)) to revise the unaligned response (to a well-aligned one) with minimum editing constraint, into a small LLM. Then, we can utilize its output revision probability for each token, to compute the reward score as:
$r_{i,j}=\begin{cases}1,&P_{re}(y_{i,j}|p_{i},x_{i},y_{i}^{+},y^{-}_{i,<j})<u\\ 0,&\text{others}\end{cases},$ (5)
where $p_{i}$ is the prompt to guide the reward model, $y_{i,j}$ is the $j$-th token in the negative response $y_{i}^{-}$, $u$ is a hyper-parameter to control the threshold. In this way, we can select the key tokens about the unalignment according to the 0-1 reward score.
Fine-grained Unlearning with NPO. Based on the selected unalignment-related key tokens, we perform unlearning to remove the unaligned knowledge in the LLM and unleash the potential of learning aligned knowledge. Concretely, we utilize the NPO method, which is the revision based on DPO and only focuses on minimizing the likelihood of generating negative responses. The objective function of NPO is as follows,
$\mathcal{L}_{NPO}(\theta)={\log{\sigma\left(-\beta\log\frac{P(y_{i}^{-}|x_{i})}{P_{\text{ref}}(y_{i}^{-}|x_{i})}\right)}}.$ (6)
Whereas, the NPO loss would also punish the tokens that are irrelevant to the unalignment but exist in the negative response. To address it, we constrain that only the key tokens are involved into loss computation, to avoid unlearning the irrelevant tokens. Formally, we decompose the objective into the token level, and add the 0-1 reward score as the token weights. Thus, the objective function can be revised as follows, and we only optimize the top-$k_{1}$most important neurons, denoted as $\mathcal{N}_{1}$,
$\mathcal{L}_{N}(\mathcal{N}_{1})=-\sum_{j=1}^{l_{i}}{\log{\sigma\left(-\beta\log\frac{P(y_{i,j}^{-}|x_{i},y_{i,<j}^{-})}{P_{\text{ref}}(y_{i,j}^{-}|x_{i},y_{i,<j}^{-})}\times r_{i,j}\right)}}.$ (7)
4.3 Alignment Improving
For the learning stage, we further improve the alignment of the LLM that has unlearned the unaligned knowledge. We adopt DPO Rafailov et al. (2023) algorithm for training, and also leverage its computed reward score to distinguish the key tokens and noisy ones.
Noisy Tokens Identification. We also identify the noisy tokens in the negative responses using the reward score in DPO, for reducing their harmful influence on learning other key tokens. As DPO requires to compare the token probabilities of the current-step LLM and its original probability, the reward of the key tokens initially own small values and increase smoothly. However, the noisy ones typically lead to large reward values, and shock the training process Chen et al. (2019). Therefore, we can utilize the reward scores dynamically computed in the DPO process, to distinguish the key and noisy tokens, denoted as:
$q_{i,j}=\begin{cases}0,&r^{\prime}_{i,j}\in\text{top }v\%\\ 1,&\text{others}\end{cases},~{}r^{\prime}_{i,j}=\frac{P(y_{i,j}^{-}|x_{i},y_{i,<j}^{-})}{P_{\text{ref}}(y_{i,j}^{-}|x_{i},y_{i,<j}^{-})},$ (8)
where $v\%$ is the hyper-parameter to control the threshold. In this way, we can identify the noisy tokens causing abnormal large rewards with weight 0, and key tokens with weight 1.
Fine-grained Learning with DPO. After obtaining the token weights, we also decompose the objective function of DPO into the token level, and add weights into the tokens from the negative response to provide fine-grained supervision. Formally, the revised objective function is as follows:
$\mathcal{L}_{D}(\mathcal{N}_{2})=-\log\sigma($ $\beta\sum_{j=1}^{l_{i}^{+}}\log\frac{P(y_{i,j}^{+}|x_{i},y_{i,<j}^{+})}{P_{\text{ref}}(y_{i,j}^{+}|x_{i},y_{i,<j}^{+})}$ (9)
-- $\beta\sum_{j=1}^{l_{i}^{-}}\log\frac{P(y_{i,j}^{-}|x_{i},y_{i,<j}^{-})}{P_{\text{ref}}(y_{i,j}^{-}|x_{i},y_{i,<j}^{-})}\times q_{i,j}),$
where we only optimize the top-$k_{2}$most important neurons, denoted as $\mathcal{N}_{2}$.
5 Experiment
5.1 Experimental Settings
Task Train / Test Dataset Number
IF Train UltraFeedback 23,976
Test AlpacaEval 2.0 805
Arena-Hard 500
QA Train ECQA 7,598
QASC 8,134
Test ECQA 2,194
QASC 926
OBQA 500
StrategyQA 687
Math Train MetaMathQA 40,000
Test GSM8k 1,319
MATH 5,000
MAWPS 2,065
TabMWP 1,000
Table 1: Statistics of the evaluation datasets. “IF” denotes the instruction following tasks.
In this section, we introduce the details of our evaluation process, including downstream datasets, baselines in the evaluation, and the implementation details of our proposed method.
Methods Question-Answering Tasks Mathematical Reasoning Tasks
ECQA QASC OBQA StrategyQA Avg. GSM8k MATH MAWPS TabMWP Avg.
SFT LLM 69.92 55.51 52.60 55.75 58.45 55.9 11.8 79.9 56.7 51.1
+ SFT 69.14 55.40 49.80 59.24 58.40 56.2 11.8 80.0 57.4 51.4
+ RFT 71.15 57.24 54.40 56.33 59.78 54.7 12.0 80.2 55.2 50.5
+ DPO 75.07 60.37 57.00 59.53 62.99 56.6 12.2 81.7 57.3 52.0
+ R-DPO 75.52 61.56 58.40 59.83 63.83 56.9 12.3 82.3 57.2 52.2
+ IPO 47.86 43.20 41.80 43.38 44.06 58.0 12.9 82.4 55.5 52.2
+ BCO 68.87 55.18 45.40 57.21 56.67 57.2 12.4 81.8 56.3 51.9
+ SimPO 62.76 52.27 46.80 53.71 53.89 57.9 12.8 82.1 56.7 52.4
+ NPO 70.56 56.59 52.80 56.04 59.00 56.4 12.3 80.1 56.5 51.3
+ Vanilla PPO 70.65 55.29 53.40 56.33 58.92 55.2 11.6 79.4 56.5 50.7
+ PPO A2C 71.06 55.18 53.00 58.37 59.40 55.2 11.7 82.1 55.8 51.2
+ ALLO 75.93 62.31 59.60 60.84 64.67 56.6 13.0 82.5 58.1 52.6
Table 2: Experimental results on question answering tasks and mathematical reasoning tasks. Avg. is the average accuracy of all sub-tasks. The best is denoted in bold and the second best is underlined.
Methods Instruction Following Tasks
AlpacalEval 2.0 Arena-Hard Avg.
SFT LLM 50.00 50.00 50.00
+ SFT 49.44 61.50 55.47
+ RFT 50.06 53.70 51.88
+ DPO 53.80 68.30 61.05
+ R-DPO 54.00 72.20 63.10
+ IPO 56.35 71.00 63.68
+ BCO 54.79 71.80 63.30
+ SimPO 54.92 69.30 62.11
+ NPO 50.06 51.10 50.58
+ Vanilla PPO 48.75 48.20 48.48
+ PPO A2C 53.50 57.80 55.65
+ ALLO 55.78 74.90 65.34
Table 3: Experimental results on instruction following tasks. Avg. is the average win rate of all sub-tasks. The best are in bold and the second-best are underlined.
Datasets. We conduct the three downstream scenarios for the comprehensive evaluation, i.e., question-answering (QA), mathematical reasoning, and instruction following. The statistics information of each task is presented in Table 1.
$\bullet$ QA tasks require LLMs to perform multi-step reasoning to solve problems. We adopt ECQA Aggarwal et al. (2021), QASC Khot et al. (2020), OpenbookQA Mihaylov et al. (2018), and StrategyQA Geva et al. (2021) as the evaluation tasks. LLMs are fine-tuned on the training set of ECQA and QASC to adapt to the QA tasks.
$\bullet$ Mathematica reasoning tasks include four challenge tasks, i.e., GSM8k Cobbe et al. (2021), MATH Hendrycks et al. (2021), MAWPS Koncel-Kedziorski et al. (2016), and TabMWP Lu et al. (2023), containing problems with different levels of difficulty. To complete the mathematical knowledge and ability of LLMs, MetaMathQA Yu et al. (2023) has been utilized to fine-tune the LLMs.
$\bullet$ Instruction following tasks assess the capacity of LLMs to follow human instructions. AlpacaEval 2.0 Li et al. (2023) and Arena-Hard Li et al. (2024) are considered as the downstream tasks. We adopt the alpaca dataset Taori et al. (2023) to fine-tune the base LLMs and UltraFeedback dataset Cui et al. (2023) for the further training process (e.g.,, DPO, ALLO).
For QA tasks and mathematical tasks, accuracy has been adopted as the evaluation metric. For the instruction following tasks, we employ gpt-3.5-turbo as the judge model and report the win rate over the backbone model (i.e., SFT LLM).
Baselines. We incorporate three categories of methods in the evaluation, including supervised fine-tuning (i.e., SFT Ouyang et al. (2022) and RFT Liu et al. (2023)), reinforcement learning (i.e., Vanilla PPO Schulman et al. (2017) and PPO A2C Mnih et al. (2016)), and alignment without RL (i.e., DPO Rafailov et al. (2023), R-DPO Park et al. (2024), IPO Azar et al. (2024), BCO Jung et al. (2024), SimPO Meng et al. (2024b), and NPO Zhang et al. (2024)).
Implementation Details. In the experiment, we fine-tune LLaMA 2 7B Touvron et al. (2023) on instruction datasets corresponding to the downstream scenarios to obtain the backbone model (i.e., SFT LLM), and conduct further training processes based on this model in the evaluation. The details of hyper-parameters are presented in Table 5.
Forgetting Stage Learning Stage QASC OBQA MATH MAWPS AlpaceEval 2.0 Arena-Hard
TLR Mask TLR Mask Acc. (%) Acc. (%) Acc. (%) Acc. (%) WR (%) WR (%)
Top-k Top-k 62.31 59.60 13.0 82.5 55.78 74.90
Top-k Top-k 62.42 59.00 12.4 82.5 55.60 75.10
Top-k Top-k 61.56 58.20 12.7 82.3 55.47 73.20
Top-k 61.77 58.80 13.1 82.4 55.22 73.20
Top-k 61.66 58.20 12.7 81.7 53.98 69.70
Last-k Top-k 62.20 59.40 12.5 82.5 55.29 70.80
Top-k Last-k 61.77 59.00 10.2 70.9 51.74 61.20
- - Top-k 62.20 59.20 12.3 82.2 55.60 72.60
Top-k - - 56.16 53.20 11.8 79.9 51.37 50.20
Table 4: The results of ablation study. “Acc.” and “WR” denote accuracy and win rate, respectively. “TLR” denotes the whether adopting token-level rewards in each stage. “Mask” indicates the neuron masking mechanism.
5.2 Main Results
The results of ALLO and baseline approaches in our evaluation are presented in Table 2 and Table 3.
According to the evaluation, we can observe that ALLO outperforms other baselines in almost all downstream scenarios and makes a great improvement over NPO and DPO, which are the backbone methods of ALLO. That is because ALLO makes great efforts to reduce the redundant elements in the alignment process, including neurons in LLMs and tokens in training data. Experimental results have shown the effectiveness of ALLO.
Besides, comparing the performance between the algorithm with fine-grained supervision signals (e.g., ALLO, PPO A2C) and the algorithm without them (e.g., DPO, Vanilla PPO), the effectiveness of the fine-grained supervision signals has been verified. Specifically, PPO A2C has achieved a 55.65% average win rate in instruction following tasks, while Vanilla PPO only achieved 48.48%. Instance-level supervision cannot focus on the details in the training data, which will optimize the erroneous parts and hurt the performance of the training methods. In contrast, token-level supervision signals can better identify whether the token is worthy to be learned, which reduces the redundancy of training content.
Moreover, the improvement brought by the unlearning method (i.e., NPO) has demonstrated that aligned and unaligned knowledge are both stored in LLMs. In the training process of NPO, the LLMs are not exposed to new knowledge and new capacities, and only are guided to forget the unaligned knowledge. This phenomenon further verifies the importance of the unlearning stage and the existence of redundant neurons in LLMs. Without the redundant neurons, is difficult of LLMs to learn both aligned and unaligned knowledge simultaneously.
Finally, we can observe that ALLO outperforms DPO and its various (e.g., R-DPO, SimPO) in all downstream scenarios, especially in the instruction following tasks. This is because DPO and its various guide LLMs to learn the positive and negative instances simultaneously, which will make LLMs confused about the aligned components in the negative instances. In contrast, ALLO first utilizes the unlearning process to lose the probability distribution in LLMs and leverage the fine-grained supervision signals to indicate the redundant tokens in the training data, to enhance the training efficiency.
Figure 3: The experimental results of the influence of different warm-up methods on downstream tasks.
5.3 Detailed Analysis
To further analyze our proposed ALLO, we conduct the ablation study, and analyze the influence of different warm-up methods and neuron mask ratio. Besides, we present a case study in Appendix D.
Ablation Study. To assess the effectiveness of each module in ALLO, we conduct the ablation study and present the evaluation results in Table 4. According to the results, we can observe that removing any component of ALLO will hurt the performance, which has verified each module in ALLO is necessary and contributes to the final results of ALLO. Besides, in QA tasks, the results of removing the neuron mask and adopting the Last-k neuron mask indicate the existence of redundant neurons in LLMs, which is the same as our empirical study. For details, even adopting the Last-k neuron mask in Stage 2 (e.g., 59.00% accuracy of OBQA) outperforms the variant without neuron masking (e.g., 58.20% accuracy of OBQA). That is because training the whole neurons in LLMs will decrease the training efficiency, and redundant updates affect the performance of downstream tasks. Moreover, without the forgetting stage, ALLO still performs better than DPO in most tasks. The reason is that the token-level reward and the neuron masking mechanism reduce the redundancy and make the training process focus on effective details in the training instances, making better utilization of the information in the dataset.
Figure 4: The experimental results of the different neuron mask ratios on ECQA and AlpaceEval 2.0, reporting the accuracy and win rate respectively. In the evaluation, we keep the mask ratio of one stage frozen and change the ratio of another stage.
Influence of Different Warm-up Methods. To assess the influence of different warm-up methods (i.e., DPO, SFT, and NPO), we conduct the relative experiment and present the results in Figure 3. In all of the evaluation tasks, leveraging DPO to warm up LLMs and select important neurons has achieved the best performance that other warm-up methods. Whether SFT or NPO, these training methods only utilize a single part of the training dataset, i.e., the positive responses or the negative responses, respectively. However, positive responses indicate the knowledge that LLMs should possess, and negative responses can locate unaligned knowledge stored in LLMs. These responses are both important and necessary in selecting the key neurons for the corresponding scenario. DPO can leverage the information in this data and guide LLMs to learn the aligned knowledge and eliminate unaligned one. In this warm-up process, the neurons related to downstream tasks will be modified largely, causing the large value of the gradient, which can more precisely locate the important neurons for the following training process.
Analysis the Ratio of Neuron Mask. We present the results of different ratios of neuron masks on the QA task (i.e., ECQA) and the instruction following task (i.e., AlpaceEval 2.0) in Figure 4. According to the evaluation results, we can observe that the performance first increases and then decreases, with the change of the neuron mask ratio. Concretely, for the ECQA task, selecting 10% neurons in the learning stage achieves the best performance, while selecting fewer or more neurons will hurt the accuracy of LLMs on downstream tasks. The increasing stage indicates that there are still several important neurons not been selected, which affects LLMs learning task-specific knowledge and abilities. After the increasing stage, the selected neurons set $\mathcal{N}$ contains more and more redundant neurons, interfering the learning process of other neurons and hurting the performance of the LLMs. The evaluation results have verified the existence of redundant updates in LLM alignment and shown that training an appropriate amount of neurons can reduce the redundancy and enhance the performance of LLMs.
6 Conclusion
In this paper, we proposed ALLO, an alignment method with low-redundant optimization, to train the most related neurons with the most useful supervised signals. In ALLO, we first estimated the importance of neurons in the LLM based on the weight changes of a reference model, and located the most related neurons for optimization. Then, we decomposed the alignment process into the forgetting and learning stages, where we leveraged token-level reward and DPO reward scores to identify the key tokens, and computing loss on them for training. Experimental results on question-answering tasks, mathematical reasoning tasks, and instruction following tasks have shown the effectiveness of ALLO.
As future work, we will consider leveraging ALLO on other important scenarios, e.g., reducing hallucination. Besides, we will also implement ALLO in larger LLMs and multimodal LLMs to validate its effectiveness.
Limitations
In this section, we discuss the limitations of our work. First, we only conduct the experiment of ALLO on 7B LLMs, with the evaluation of the LLMs with larger scaling of parameters, because of the limitation of computation resources. Actually, we comprehensively assess the performance of ALLO and the existing competitive baseline methods in various downstream tasks, and the experiment results have verified the effectiveness of our proposed methods. Second, we adopt complex reasoning and human alignment tasks in our evaluation, which mainly assess the helpfulness of LLMs. The performance of ALLO on other aspects, e.g., reducing hallucination and generating harmless response, has not been verified in this work. We leave it as future work. Finally, we do not consider the potential risk of ethics risk during LLM deployment and will investigate this issue in the future.
Appendix A Algorithm of ALLO
We present the pipeline of ALLO in Algorithm 1. The process of ALLO includes three stages, i.e., locating key neurons, unaligned knowledge forgetting, and alignment improving.
Appendix B Details of Hyper-Parameters
To better understand and reproduce our proposed ALLO, we presented the hyper-parameters in ALLO in Table 5. The hyper-parameters are a little different between different downstream tasks, that is because these tasks are in different difficulty levels and require different abilities of LLMs. It should be noted that, to conduct a fair comparison, the hyper-parameters of baseline methods are also adjusted to adapt to the corresponding tasks for better performance.
Appendix C Prompt Templates of ALLO
In ALLO, we utilize prompts to guide the teacher model to rewrite the generated response from student models and induce the student model to solve the downstream tasks. The templates of the prompt in ALLO are presented in Table 6. For the solution rewriting process, we feed the problem, ground-truth reference, and generated response into the teacher model, with the instruction of rewriting in the prefix. Besides, for the downstream tasks, the instruction prefix and problem will be given into LLMs.
Appendix D Case Study
To better demonstrate our proposed ALLO, we present the case study on QA task (i.e., ECQA) in Table 7. In this case, we can observe that LLM after DPO training still cannot catch the relation between “tickets” and the destination John needed to go to, and focus on the relation between “cross country” and “race track”. This phenomenon has shown that unaligned knowledge is not eliminated and still exists in LLMs after DPO training. In contrast, after ALLO training, LLM can correctly seize on the key elements of the problem (i.e., “ticket”) and perform reasoning along the correct direction. That is because low-redundant optimization can reduce the redundant updates in the alignment process and make LLMs focus on the key knowledge and information.

### v2

@@
-Low-Redundant Optimization for Large Language Model Alignment
+Not Everything is All You Need: Toward Low-Redundant Optimization for Large Language Model Alignment
 Abstract
@@
 The key issue is that LLMs might overfit into the unexpected patterns or superficial styles in the human preference data (i.e., pairs of positive-negative response) Du et al.
-(2024).
+(2024a).
 It is the side effect of their powerful learning capability derived from the large-scale trainable parameters Song et al.
@@
 (2023); Zhou et al.
-(2024).
+(2024b).
 However, given the large expenses of the LLM training, existing work Hu et al.
@@
 Following the unlearning mechanism, in this work, we utilize an unlearning algorithm to correct the unaligned knowledge stored in the neurons of LLMs.
+Low-Cost LLM Training.
+Due to the high computation cost of LLM training, previous work either trains a small portion of parameters in LLMs, or trains small-scale LLMs.
+For the former, existing work has shown the parameter redundancy in LLMs, and proposed related approaches to reduce the trainable parameters in LLM Hu et al.
+(2022); Li and Liang (2021).
+Besides, quantization has also been used to further reduce the requirements of GPU memory Dettmers et al.
+(2023).
+Moreover, recent work has further improved the above methods by adjusting the updated parameters within LLMs or improving the training strategies Zhou et al.
+(2024a); Du et al.
+(2024b).
+For training small LLMs, to break the ceiling derived from the scaling law, knowledge distillation methods have been widely used for transferring the capabilities from the teacher model to small LLMs He et al.
+(2023).
+However, it is also promising to small models for hyper-parameters searching, to obtain the optimal training hyper-parameters for large LLMs Hu et al.
+(2024).
+In this work, we mainly focus on training a small portion of parameters to reduce the useless redundancy in the human alignment training process.
+Such a way is capable of reducing the training cost and also improves the performance.
 3 Preliminary
@@
 Finally, we do not consider the potential risk of ethics risk during LLM deployment and will investigate this issue in the future.
+Acknowledgement
+This work was partially supported by National Natural Science Foundation of China under Grant No.
+62222215, Beijing Natural Science Foundation under Grant No.
+L233008 and 4222027.
+Xin Zhao is the corresponding author.
 Appendix A Algorithm of ALLO